# CREA: A Collaborative Multi-Agent Framework for Creative Content Generation with Diffusion Models

### v1

@@ 4 Experiments
Experimental Setup In this section, we evaluate our method’s ability to generate highly creative edits and images. All experiments use FLUX.1-dev111https://huggingface.co/black-forest-labs/FLUX.1-dev. For editing, we employ ControlNet222https://huggingface.co/InstantX/FLUX.1-dev-Controlnet-Union with a conditioning scale of 0.4. For image generation, we vary the CFG scale from 3.5 to 40 to explore different levels of creativity and control. Additionally, we utilize Autogen [72] for our agentic framework. All experiments are performed on an NVIDIA L40 GPU. We use GPT-4o as our MLLM for all agents [51]. We use QWEN2.5-VL-32B for MLLM as a judge metrics [15]. We set the Creativity Index () to 24 for editing tasks and 26 for generation tasks with maximum number of iterations as . A full run of the pipeline takes approximately 3–5 minutes, depending on the number of self-enhancement rounds. Our source code will be publicly available.
@@ 4.1 Qualitative Results
Creative Image Editing First, we qualitatively showcase how CREA transforms input images into various creative modifications. As seen in Fig. 3, CREA’s autonomous agents analyze the input image or concept and generate creative prompts—e.g., focusing on style, color, or thematic twists—without requiring extensive user intervention. Since no existing method is explicitly designed for creative image editing, we adapt state-of-the-art editing models for comparison. Specifically, we evaluate the following baselines: LEDITS++ [6], InstructPix2Pix [7], SDEdit [46] and TurboEdit [11]. For a fair comparison, we apply a “creative <object>” prompt (e.g., “a creative couch”) to the baseline methods, mirroring the objective of CREA. As shown in Fig. 4, standard editing methods often fail to generate distinctly creative concepts: InstructPix2Pix typically adds vibrant colors or alters the background extensively without fundamentally reimagining the object. LEDITS++, SDEdit, and TurboEdit struggle to introduce creative features beyond superficial stylistic changes. In contrast, CREA successfully performs creative edits in a disentangled manner.
@@
Creative Image Generation For creative image generation, we compare CREA against ConceptLab [55], which is the closest related approach, as well as two generative baselines: Flux [14] and SDXL [53]. To ensure a fair comparison, we use the same random seed for each method in every evaluated prompt, allowing for direct visual comparisons. As illustrated in Fig. 5, CREA consistently produces diverse and creative generations across multiple domains. In contrast, ConceptLab struggles to maintain the intended concept, particularly for highly abstract or unconventional categories. For example, when generating “a monster,” ConceptLab often fails to produce a meaningful interpretation, as it relies on extracting subcategories from the BLIP model. If a given concept lacks well-defined subcategories or consists of highly specific attributes, ConceptLab’s fails to perform well. CREA, on the other hand, generalizes across a broad range of creative categories without relying on predefined subcategory constraints.

### v2

@@ 4 Experiments
Experimental Setup In this section, we evaluate our method’s ability to generate highly creative edits and images. All experiments use FLUX.1-dev [flux2024]. For editing, we employ ControlNet [zhang2023adding] with a conditioning scale of 0.4 and Canny as the condition. For image generation, we vary the CFG scale from 3.5 to 40 to explore different levels of creativity and control. Additionally, we utilize Autogen [wu2024autogen] for our agentic framework. All experiments are performed on a 48GB NVIDIA L40 GPU. We use GPT-4o as our MLLM for all agents [achiam2023gpt]. We set the Creativity Index () threshold to 24 for editing, requiring a majority of criteria to score at least 4 (“very good”) for an image to be considered creative. For generation, we use a higher threshold of 26 to reflect greater creative freedom unconstrained by a base image. We cap the number of refinement iterations at . A full run of the pipeline takes approximately 3–5 minutes, depending on the number of self-enhancement rounds. Our source code is publicly available at https://crea-diffusion.github.io.
@@ 4.1 Qualitative Results
Creative Image Editing First, we qualitatively showcase how CREA transforms input images into various creative modifications. As seen in Fig. 3, CREA’s autonomous agents analyze the input image or concept and generate creative prompts: e.g., focusing on style, color, or thematic twists, without requiring extensive user intervention. Since no existing method is explicitly designed for creative image editing, we adapt state-of-the-art editing models for comparison. Specifically, we evaluate the following baselines: LEDITS++ [brack2024ledits++], InstructPix2Pix [brooks2023instructpix2pix], SDEdit [meng2021sdedit] and TurboEdit [deutch2024turboedit]. For a fair comparison, we apply a “creative <object>" prompt (e.g., “a creative couch") to the baseline methods, mirroring the objective of CREA. As shown in Fig. 4, standard editing methods often fail to generate distinctly creative concepts: InstructPix2Pix typically adds vibrant colors or alters the background extensively without fundamentally reimagining the object. LEDITS++, SDEdit, and TurboEdit struggle to introduce creative features beyond superficial stylistic changes. In contrast, CREA successfully performs creative edits in a disentangled manner. For more experiments such as more fine-grained editing examples, or other image generators, please refer to Appendix D.4.
@@
Creative Image Generation For creative image generation, we compare CREA against ConceptLab [richardson2024conceptlab], which is the closest related approach, as well as two generative baselines: Flux [flux2024] and SDXL [podell2023sdxl]. To ensure a fair comparison, we use the same random seed for each method in every evaluated prompt, allowing for direct visual comparisons. As illustrated in Fig. 5, CREA consistently produces diverse and creative generations across multiple domains. In contrast, ConceptLab struggles to maintain the intended concept, particularly for highly abstract or unconventional categories. For example, when generating “a monster," ConceptLab often fails to produce a meaningful interpretation, as it relies on extracting subcategories from the BLIP model. If a given concept lacks well-defined subcategories or consists of highly specific attributes, ConceptLab’s fails to perform well. CREA generalizes across a broad range of creative categories without relying on predefined subcategories.
4.2 Quantitative Results
Creative Image Editing. For image editing, we compare our method against state-of-the-art techniques, including LEDITS++, InstructPix2Pix, SDEdit, and TurboEdit, across a range of evaluation metrics. To assess how well the edited objects align with the text descriptions, we compute the CLIP score by measuring the similarity between the generated image and its corresponding text prompt. For evaluating diversity, we use the LPIPS [zhang2018unreasonable] score, which quantifies perceptual distance between images. Unlike conventional usage where lower LPIPS indicates better reconstruction, a higher LPIPS in our case signifies stronger, more transformative edits, a desirable trait in creative tasks. Additionally, we compute the DINO [zhang2022dino] score which measures how well the edited image retains key structural and semantic characteristics of the original image, and Vendi [friedman2022vendi] score, which quantifies diversity by calculating the Shannon entropy of the eigenvalues of the similarity matrix among the generated images. As shown in Table 1, our method consistently outperforms all baselines across all evaluation metrics, demonstrating its superior ability to generate creative and diverse image edits.
Creative Image Generation. For creative image generation, we compare CREA against ConceptLab, Flux, and SDXL. As shown in Table 1, our method achieves superior results in LPIPS and VENDI scores, while in CLIP scores, it was outperformed by SDXL. We observe that SDXL tends to generate highly colorful images, which aligns with CLIP’s depiction of creativity. However, it produces the lowest LPIPS scores, indicating that its generated images are highly similar to each other, and it also achieves significantly lower VENDI scores compared to CREA and ConceptLab. While ConceptLab and our method achieve comparable results overall, ConceptLab has significantly lower performance CLIP scores, suggesting that its generated images do not align well with the text prompt. This limitation stems from ConceptLab’s reliance on well-defined subcategories—a core design principle of their method. As shown in Fig. 5, ConceptLab fails when a concept lacks sufficient subcategories, such as its inability to accurately generate a couch object. While ConceptLab attains a high VENDI score, this failure highlights its trade-off between diversity and semantic alignment.
For additional comparisons with recent methods such as GenArtist [wang2024genartist] and RF-Inversion [rout2024semantic], see Appendix D.3.
User Study. To assess the creativity of the generated images from a human perception standpoint, we conducted a user study on Prolific.com with 50 participants (see Table 1. Participants rated each criterion using a 5-point Likert scale, ranging from 1 (Not at all) to 5 (Very well). For the image generation task, User Study-Q1 evaluates Usability, measuring how accurately the generated image represents the specified object. User Study-Q2 evaluates Creativity, assessing the uniqueness and originality of the image. Among all competitors, SDXL performed the best on Q1 since it is generating generic objects (e.g., a simple cup) which is perceived as more useful by participants, whereas our method significantly outperformed others on Q2 due to its ability to produce more creative visuals. For edited images, we evaluate editing consistency—how well the original image characteristics are preserved while performing edits (User Study-Q1)—and the creativity of the edits (User Study-Q2). Our method achieves comparable results to state-of-the-art editing techniques on Q1, demonstrating strong disentangled editing capabilities. Moreover, it surpasses all competitors on Q2, highlighting its superior ability to generate creative edits.
4.3 Additional Experiments
User-Guided Editing and Generation Our method enables users to guide the creative process according to their preferences (see Fig. 6 (a)). For example, users can specify not only that a creative edit should be applied to a cup but also include additional conditions, such as ‘monster’, to influence the generation. Our approach incorporates these preferences, steering the creative process accordingly. Compared to the baseline Flux + ControlNet approach, where we provided the prompt ‘A creative <condition> <object>’, our results demonstrate greater creativity.
Personalization We also present personalization results using CREA-generated creative images using an off-the-shelf personalization adapter [wang2024instantstyle]. Fig. 6 (b) demonstrates how generated subjects can be adapted to various contexts or how their styles can be transferred to create new images. Note that due to the limitations in personalization models, some fine-grained details were not able to capture by [wang2024instantstyle], however the main characteristics are still preserved.
Video Generation To demonstrate the versatility and extensibility of our agentic framework, we extend CREA to creative video generation, where an initial prompt (e.g., ‘a train’) is automatically transformed into creative scenes. We conduct experiments using the CogVideoX [yang2024cogvideox] model as the generative backbone. Instead of initializing the process with a static creative blueprint, our Creative Director agent generates a structured creative video plan containing key fields: Subject, Action, Setting, Style, and optional Additional Details. This plan is passed to the Prompt Architect, who composes a coherent and high-creativity video prompt via contrastive prompt fusion, similar to our image generation pipeline. The Generative Executor then synthesizes the video using CogVideoX. This evaluation serves as a proof of concept for applying our multi-agent creativity principles beyond static imagery. As shown in Figure 7, our method produces significantly more imaginative and visually engaging results compared to baseline model.
| Method | LPIPS-Diversity | Vendi |
| Base | 0.302 0.129 | 3.19 1.49 |
| +Principles | 0.312 0.142 | 3.59 2.46 |
| +Contrastive | 0.391 0.150 | 3.50 1.64 |
| +Self-Enhancement | 0.414 0.157 | 3.70 1.97 |
Ablation Studies We perform comprehensive ablation studies to analyze the design choices of CREA (see Fig. 8). (a) Ablation on Generalization to other T2I models: Our method extends effectively to different generative models, such as SDXL and DALL-E. (b) Ablation on Parameter Sensitivity: We ablate CFG values (3.5-40) for Flux and the conditioning scale for ControlNet (0.1-0.5) to evaluate their impact. (c) Ablation on Iterative Refinement: We demonstrate the benefits of our method’s refinement process over multiple iterations. (d) Ablation on Prompt Variations: We explore alternative prompts beyond ‘a creative <obj>’ to steer the model away from conventional generations. (e) Ablation on Model Components We also conduct an ablation study on our key components (see Table 2). The base version represents a baseline where the multi-modal LLM model is simply prompted to generate a creative description for a given object. Please see Appendix for more ablations, such as an ablation with negative prompting.
5 Discussion
Limitations and Broader Impact Since our framework utilizes a T2I model such as Flux, certain biases inherent to the generative backbone can affect the outputs. For instance, we observe cases where the background is unintentionally darkened based on the object’s semantics (e.g., editing a “lighted box” often results in darker surroundings to emphasize contrast), or where the model hallucinates human figures even when not prompted (see Fig. 8 (e)). Moreover, we recognize that defining and quantifying visual creativity is itself an open research problem. Instead of introducing a new assessment protocol, we tackle creative image editing and generation through established principles in creativity. Assessing what is truly a creative output, however, remains beyond the scope of this paper. For broader impacts, on the positive side, our agentic framework can democratize high-level creative tools, lowering barriers for artists, educators, and hobbyists to experiment with sophisticated visual storytelling. Potential downsides include possible displacement of human labor in creative industries. We encourage responsible deployment by fostering human-in-the-loop usage that positions the system as a co-creative assistant rather than a replacement for human creators.
Conclusion In this work, we introduce a novel agentic framework for creative image editing and generation, pioneering a disentangled approach that enables greater flexibility and artistic control. By leveraging specialized agents that collaborate to refine and enhance outputs, our method overcomes the limitations of traditional prompt-to-image models, reducing the burden on users while fostering creativity. We demonstrate the versatility of our approach across editing and generation, and highlight its potential for creative video generation. Our findings suggest that agentic frameworks can serve as a powerful foundation for more autonomous and creative AI systems, opening new directions for creative collaboration between humans and AI.
Appendix A Investigating Creativity Types
We investigate emerging patterns in generated images to identify key attributes—shape, color, and texture that influence creativity.Our analysis reveals that our method achieves significantly higher scores in these aspects, leading to the generation of more visually diverse and imaginative objects (see Fig.9).Compared to baseline models, our approach demonstrates a stronger ability to manipulate these factors, reinforcing its effectiveness in producing unique and aesthetically rich generations.